# TempFlow-GRPO: When Timing Matters for GRPO in Flow Models

### v1

@@ Abstract
Recent flow matching models for text-to-image generation have achieved remarkable quality, yet their integration with reinforcement learning for human preference alignment remains suboptimal, hindering fine-grained reward-based optimization. We observe that the key impediment to effective GRPO training of flow models is the temporal uniformity assumption in existing approaches: sparse terminal rewards with uniform credit assignment fail to capture the varying criticality of decisions across generation timesteps, resulting in inefficient exploration and suboptimal convergence. To remedy this shortcoming, we introduce TempFlow-GRPO (Temporal Flow GRPO), a principled GRPO framework that captures and exploits the temporal structure inherent in flow-based generation. TempFlow-GRPO introduces two key innovations: (i) a trajectory branching mechanism that provides process rewards by concentrating stochasticity at designated branching points, enabling precise credit assignment without requiring specialized intermediate reward models; and (ii) a noise-aware weighting scheme that modulates policy optimization according to the intrinsic exploration potential of each timestep, prioritizing learning during high-impact early stages while ensuring stable refinement in later phases. These innovations endow the model with temporally-aware optimization that respects the underlying generative dynamics, leading to state-of-the-art performance in human preference alignment and standard text-to-image benchmarks.
@@ 1 Introduction
The majority of existing approaches Gu et al. (2024); Hong et al. (2024), including recent works like Flow-GRPO Liu et al. (2025) and DanceGRPO Xue et al. (2025), treat the multi-step generation process as a ”black box” with temporally agnostic optimization. They apply uniform updates across all timesteps despite the fact that each timestep operates under different noise conditions and contributes differently to final image quality. Specifically, we plot the Fig. 1 (left) with applying SDE at only one timestep in the entire ODE trajectory, which ensures that any deviation in the final reward can be attributed to the stochastic exploration introduced at that specific step. As shown in Fig. 1 (left), the std of the reward varies dramatically between timesteps, reaching a peak during early structural decisions (steps 0-2) and approaching zero during final refinements (steps 6-8). Yet Flow-GRPO maintains uniform treatment throughout, squandering high-impact exploration opportunities while focusing on later steps. Alternative approaches like SPO Liang et al. (2025) attempt to address temporal dynamics through process reward models, but training such models on semantically ambiguous intermediate states is notoriously difficult. This raises a fundamental question: how can we effectively achieve precise credit assignment for intermediate actions while adapting optimization intensity to each timestep’s exploration capacity?
@@
Figure 1: (Left) Reward Variance Analysis: We plot the standard deviation of PickScore at each denoising step for 200 prompts, per prompt group size is 24. The results, obtained via applying SDE at only one step, reveal that reward variance is highest in the initial steps, indicating that early-stage interventions are most impactful for exploration. (Right) Method Illustration: By branching a stochastic (SDE) exploration from a specific, known state on a deterministic (ODE) trajectory, we create a controlled experiment. The resulting difference in the final reward can be unambiguously attributed to the exploration action taken at that precise branching point.
@@
We address these limitations with TempFlow-GRPO, a temporally-aware RL framework built on two key insights. First, we define the visualization method in the left of Fig. 1 as trajectory branching, which enables precise credit assignment by strategically introducing stochasticity at individual timesteps while maintaining deterministic evolution elsewhere (Fig. 1, Right). This provides provable guarantees: (1) reward variance localizes to the branching point, (2) improvements are directly attributable to specific exploration outcomes, and (3) existing reward models require no modification. Second, noise-aware policy weighting modulates optimization intensity based on each timestep’s intrinsic noise level. Early high-noise stages receive larger weight updates to encourage structural exploration, while late low-noise stages receive gentler updates to preserve learned features. Together, these mechanisms create a framework that is conceptually simple, computationally efficient, and seamlessly integrates into existing flow matching architectures—all while respecting the temporal dynamics that uniform approaches ignore.
@@
  • We introduce trajectory branching and noise-aware weighting to learn temporally-structured policies that respect generative dynamics.
@@ 2 Related Work
Alignment for Diffusion Models. Alignment for diffusion models has become a rapidly emerging topic in generative modeling research. D3PO Yang et al. (2024) introduce the Direct Preference for Denoising Diffusion Policy Optimization method to directly fine-tune diffusion models. Diffusion-DPO Wallace et al. (2024) is adapted from the Direct Preference Optimization (DPO Rafailov et al. (2023)), a simpler alternative to RLHF which directly optimizes a policy that best satisfies human preferences under a classification objective. DyMO Xie & Gong (2025) propose a plug-and-play training-free alignment method for aligning the generated images and human preferences during inference. Recently, Flow-GRPO Liu et al. (2025), the first method to integrate online reinforcement learning (RL) into flow matching models. However, Flow-GRPO applies uniform optimization pressure across all timesteps and suffers from sparse terminal reward problems, failing to account for the time-varying exploration potential inherent in the stochastic diffusion process. Our TempFlow-GRPO addresses these limitations through trajectory branching for precise credit assignment and noise-aware policy weighting that aligns optimization pressure with the natural exploration capacity at each timestep.
Figure 2: (Left) Performance comparison on the PickScore benchmark. To ensure a fair and robust evaluation, we first present Flow-GRPO (Fixed), an improved baseline where we replace the original global std stabilization with a more appropriate group-wise std. On top of this stronger baseline, our core innovation, trajectory branching (+ Traj. Branch), delivers a significant performance leap. Our full model, TempFlow-GRPO, integrates all components to achieve the highest performance. (Right) TempFlow-GRPO’s dominance is further confirmed on the Geneval benchmark, where it substantially surpasses not only the Flow-GRPO baseline but also leading state-of-the-art models, including GPT-4o, SD3.5-M, and FLUX.
@@
Process Reward. Recent studies have demonstrated that shaping the reward process, rather than solely relying on sparse terminal rewards, can significantly accelerate learning and improve policy performance. Zhang et al. (2025) introduces a more comprehensive evaluation framework, which combines response-level and step-level metrics. ThinkPRM Khalifa et al. (2025) builds data-efficient PRMs as verbalized step-wise reward models that verify every step in the solution by generating a verification chain-of-thought (CoT). In diffusion models, SPO Liang et al. (2024) trains a separate step-aware preference model that can be applied to both noisy and clean images. Despite this, PRMs require step-level supervision, making them expensive to train. There are several methods using outcome reward to replace process reward. PRIME Cui et al. (2025) enables online PRM updates using only policy rollouts and outcome labels through implict process rewards. Math-Shepherd Wang et al. (2023) assigns a reward score to each step of math problem solutions. Due to the significant challenges in scoring noisy images, there is an urgent need for an algorithm that enables process reward on flow models. Our TempFlow-GRPO elegantly circumvents the need for specialized process reward models by directly attributing outcome-based reward signals to intermediate exploratory actions, enabling precise credit assignment without the computational overhead of training step-level evaluators for semantically ambiguous noisy states.
@@ 3 Preliminary: Flow-GRPO
Flow-GRPO enhances flow models using online RL. First, we revisit the core idea of GRPO. Then, we show how Flow-GRPO converts the deterministic ODE sampler into a SDE sampler with same marginal distribution, which introduces the stochasticity needed for applying GRPO.
@@
Convert ODE to SDE. GRPO relies on stochastic sampling to generate diverse trajectories for advantage estimation and exploration. However, flow matching models use a deterministic ODE for the forward process:
@@
Flow-GRPO converts the deterministic ODE into an equivalent SDE that matches the original model’s marginal probability density function at all timesteps. The final update rule is as follows:
@@ 4.1 Temporal Flow-GRPO
While Flow-GRPO represents a significant advancement in applying online RL to flow matching models, its direct adoption of the standard GRPO framework overlooks the unique, time-dependent dynamics inherent in the generative process. We identify two fundamental limitations that motivate our approach:
Sparse Terminal Reward Problem: The current approach relies on sparse terminal rewards, assigning the same credit uniformly across all timesteps. This fails to differentiate the critical, high-impact decisions of early generation from the fine-tuning adjustments of later stages.
Uniform Optimization Weighting: The GRPO objective applies uniform optimization among all steps, ignoring that the underlying SDE sampler possesses non-uniform noise. The potential for meaningful exploration is greater in high-noise early steps than in the low-noise refinement phase Xie & Gong (2025).
To address these limitations, as illustrated in Fig. 3, we propose TempFlow-GRPO, a unified framework that introduces two synergistic innovations: process rewards via trajectory branching and noise-aware policy weighting.
@@ 4.1.1 Trajectory Branching for Process Rewards
Consider a generative process parameterized by $\theta$. In flow matching models, the policy gradient can be written as:
Definition 1 (Trajectory Branching): We define a trajectory branching operation where a trajectory evolves deterministically at a designated branching timestep $k$, where $\bm{x}_{k}$ is sampled from initial noise $\bm{x}_{T}$ using Eq. 4. At this branching point, stochasticity is introduced via noise variable $\bm{\epsilon}$ in Eq. 5, yielding $\bm{x}_{k-1}=\text{SDE}(\bm{x}_{k},\bm{\epsilon})$. The remainder of the trajectory $\bm{x}_{k-2},\bm{x}_{k-3},\ldots,\bm{x}_{0}$ is generated deterministically as $\bm{x}_{0}=\text{ODE}(\bm{x}_{k-1})$.
@@
In practice, we replace the reward for the $k$-th step from $R(\bm{x}_{0}^{i},\bm{c})$ to $R(\text{ODE}(\text{SDE}(\bm{x}_{k}^{i},\bm{\epsilon}^{i})),\bm{c})$, where $\bm{x}_{0}^{i}$ is sampled with SDE and $\text{ODE}(\text{SDE}(\bm{x}_{k}^{i},\bm{\epsilon}^{i})$ is sampled with ODE-SDE-ODE. Trajectory branching allows for the precise attribution of the terminal reward to step $k$, effectively creating a temporally-aware reward signal.
Figure 4: (Left) A comparative analysis between the reward standard deviation (Reward Std) and the noise level over the generative steps. The two curves show a strong correlation. (Right) Scale terms reveal a mismatch in standard GRPO: Scale terms are inversely proportional to noise level, causing low-noise refinement steps to dominate optimization despite having minimal impact on image content. Early steps that establish global structure receive weak gradients, while late steps that only adjust local details produce strong gradients. Our noise-aware reweighting compensates for this inverse relationship, ensuring that optimization intensity aligns with each timestep’s actual capacity to influence the final image.
@@ 4.1.2 Noise-Aware Policy Weighting
Reweighting by Reward Std. An intuitive approach to reweighting the policy loss is to utilize the standard deviation of the rewards. As illustrated in the left of Fig. 4, the reward standard deviation exhibits a similar decaying trend as the denoising process unfolds, suggesting it could serve as a meaningful indicator of learning importance. However, this approach introduces significant complexity. It would necessitate maintaining a dynamic weighting hyperparameter for each timestep, likely tracked via an Exponential Moving Average (EMA Morales-Brotons et al. (2024)). This dynamic estimation process can be unstable and adds an undesirable layer of complexity to the training regime. Consequently, we seek a more direct and elegant weighting scheme.
Reweighting by Noise Level. We wonder if directly leverage the level of the exploration space itself as a proxy for a reweighting factor. We visualized the noise level alongside the reward standard deviation (Fig. 4, left) and observed a striking correspondence between the two. This strong correlation suggests that the noise level, serves as an excellent and intrinsic proxy for the exploration capacity and associated risk at each timestep.

### v2

@@ Abstract
Recent flow matching models for text-to-image generation have achieved remarkable quality, yet their integration with reinforcement learning for human preference alignment remains suboptimal, hindering fine-grained reward-based optimization. We observe that the key impediment to effective GRPO training of flow models is the temporal uniformity assumption in existing approaches: sparse terminal rewards with uniform credit assignment fail to capture the varying criticality of decisions across generation timesteps, resulting in inefficient exploration and suboptimal convergence. To remedy this shortcoming, we introduce TempFlow-GRPO (Temporal Flow GRPO), a principled GRPO framework that captures and exploits the temporal structure inherent in flow-based generation. TempFlow-GRPO introduces three key innovations: (i) a trajectory branching mechanism that provides process rewards by concentrating stochasticity at designated branching points, enabling precise credit assignment without requiring specialized intermediate reward models; (ii) a noise-aware weighting scheme that modulates policy optimization according to the intrinsic exploration potential of each timestep, prioritizing learning during high-impact early stages while ensuring stable refinement in later phases; and (iii) a seed group strategy that controls for initialization effects to isolate exploration contributions. These innovations endow the model with temporally-aware optimization that respects the underlying generative dynamics, leading to state-of-the-art performance in human preference alignment and text-to-image benchmarks.
Figure 1: Images generated by our proposed TempFlow-GRPO with FLUX.1-dev. It substantially improves the baseline models, achieving superior photorealism and enhanced fine-grained detail.
@@ 1 Introduction
The majority of existing approaches Gu et al. (2024); Hong et al. (2024), including recent works like Flow-GRPO Liu et al. (2025) and DanceGRPO Xue et al. (2025), treat the multi-step generation process as a ”black box” with temporally agnostic optimization. They apply uniform updates across all timesteps despite the fact that each timestep operates under different noise conditions and contributes differently to final image quality. Specifically, we plot the Figure 2 (left) with applying SDE at only one timestep in the entire ODE trajectory, which ensures that any deviation in the final reward can be attributed to the stochastic exploration introduced at that specific step. As shown in Figure 2 (left), the std of the reward varies dramatically between timesteps, reaching a peak during early structural decisions (steps 0-2) and approaching zero during final refinements (steps 6-8). Yet Flow-GRPO maintains uniform treatment throughout, squandering high-impact exploration opportunities. Alternative approaches like SPO Liang et al. (2025) attempt to address temporal dynamics through process reward models, but training such models on semantically ambiguous intermediate states is notoriously difficult. This raises a fundamental question: how can we effectively achieve precise credit assignment for intermediate actions while adapting optimization intensity to each timestep’s exploration capacity?
@@
Figure 2: (Left) Reward Variance Analysis: We plot the standard deviation of PickScore at each denoising step for 200 prompts, per prompt group size is 24. The results, obtained via applying SDE at only one step, reveal that reward variance is highest in the initial steps, indicating that early-stage interventions are most impactful for exploration. (Right) Method Illustration: By branching a stochastic (SDE) exploration from a specific, known state on a deterministic (ODE) trajectory, the resulting difference in the final reward can be unambiguously attributed to the exploration action taken at that precise branching point.
@@
We address these limitations with TempFlow-GRPO, a temporally-aware RL framework built on two key insights. First, we define the visualization method in the left of Figure 2 as trajectory branching, which enables precise credit assignment by strategically introducing stochasticity at individual timesteps while maintaining deterministic evolution elsewhere (Figure 2, Right). This provides provable guarantees: (1) reward variance localizes to the branching point, (2) improvements are directly attributable to specific exploration outcomes, and (3) existing reward models require no modification. Second, noise-aware policy weighting modulates optimization intensity based on each timestep’s intrinsic noise level. Early high-noise stages receive larger weight updates to encourage structural exploration, while late low-noise stages receive gentler updates to preserve learned features. Third, we introduce a seed-level grouping strategy that controls for initial noise influence, ensuring reward variations are attributed to branching exploration rather than random initialization. Together, these mechanisms create a framework that is conceptually simple, computationally efficient, and seamlessly integrates into existing flow matching architectures—all while respecting the temporal dynamics that uniform approaches ignore.
@@
  • We introduce trajectory branching and noise-aware reweighting mechanisms to learn temporally-structured policies that respect the inherent dynamics of generative models. Additionally, we propose an efficient seed group strategy that effectively isolates exploration effects and considerably enhances overall performance.
@@ 2 Related Work
Alignment for Diffusion Models. Alignment for diffusion models has emerged as a key research area. D3PO Yang et al. (2024) introduces Direct Preference for Denoising Diffusion Policy Optimization to directly fine-tune diffusion models. Diffusion-DPO Wallace et al. (2024) adapts Direct Preference Optimization (DPO Rafailov et al. (2023)) as a simpler RLHF alternative that directly optimizes policies satisfying human preferences. DyMO Xie & Gong (2025) proposes a training-free alignment method for inference-time preference alignment. Flow-GRPO Liu et al. (2025) first integrates online reinforcement learning into flow matching models. However, Flow-GRPO applies uniform optimization across timesteps and suffers from sparse terminal rewards, ignoring time-varying exploration potential in stochastic processes. Our TempFlow-GRPO addresses these limitations through trajectory branching for precise credit assignment and noise-aware policy weighting aligned with natural exploration capacity.
Figure 4: (Left) Performance comparison on the PickScore benchmark across training steps and GPU hours. Flow-GRPO (Prompt) represents an improved baseline with group-wise standard deviation stabilization. TempFlow-GRPO consistently outperforms both Flow-GRPO variants in both sample efficiency (steps) and computational efficiency (GPU hours), demonstrating superior training efficiency while achieving the best performance. (Right) On the Geneval benchmark, TempFlow-GRPO achieves the highest performance, significantly outperforming Flow-GRPO and surpassing state-of-the-art models including GPT-4o, FLUX.1 Dev, and SD3.5-Medium.
@@
Process Reward. Shaping reward processes beyond sparse terminal rewards significantly improves learning and policy performance. Zhang et al. Zhang et al. (2025) combines response-level and step-level evaluation metrics. ThinkPRM Khalifa et al. (2025) builds data-efficient process reward models (PRMs) using verification chain-of-thought for step-wise verification. In diffusion models, SPO Liang et al. (2024) trains step-aware preference models for both noisy and clean images. However, PRMs require expensive step-level supervision. Methods like PRIME Cui et al. (2025) enable online PRM updates using only outcome labels through implicit process rewards. Math-Shepherd Wang et al. (2023) assigns reward scores to solution steps. Given challenges in scoring noisy images, algorithms enabling process rewards for flow models are needed. Our TempFlow-GRPO circumvents specialized process reward models by directly attributing outcome-based signals to intermediate actions, enabling precise credit assignment without computational overhead of training step-level evaluators for semantically ambiguous states.
@@ 3 Preliminary: Flow-GRPO
Convert ODE to SDE. Flow-GRPO converts the deterministic ODE into an equivalent SDE that matches the original model’s marginal probability density function at all timesteps. The ODE and SDE is as follows:
@@ 4.1 Temporal Flow-GRPO
Flow-GRPO advances online RL for flow matching but overlooks time-dependent generative dynamics. We identify two key limitations: (1) Sparse Terminal Rewards: uniform credit assignment across timesteps fails to distinguish critical early decisions from later fine-tuning; (2) Uniform Optimization: ignoring non-uniform noise in SDE sampling, where high-noise early steps offer greater exploration potential than low-noise refinement phases Xie & Gong (2025). To address these issues, we propose TempFlow-GRPO (Figure 5), introducing process rewards via trajectory branching and noise-aware policy weighting.
@@ 4.1.1 Trajectory Branching for Process Rewards
Definition 1 (Trajectory Branching): We define a trajectory branching operation where a trajectory evolves deterministically until a designated branching timestep $k$, where $\bm{x}_{k}$ is obtained from initial noise $\bm{x}_{T}$ using Equation 4. At this branching point, stochasticity is introduced via noise variable $\bm{\epsilon}$ in Equation 5, yielding $\bm{x}_{k-1}=\text{SDE}(\bm{x}_{k},\bm{\epsilon})$. The remainder of the trajectory $\bm{x}_{k-2},\bm{x}_{k-3},\ldots,\bm{x}_{0}$ is generated deterministically as $\bm{x}_{0}=\text{ODE}^{k-1}(\bm{x}_{k-1})$, $\text{ODE}^{k-1}$ denotes k1k-1 times ODE.
@@
In practice, we replace the reward for the $k$-th step from $R(\bm{x}_{0}^{i},\bm{c})$ to $R(\text{ODE}^{k-1}(\text{SDE}(\bm{x}_{k}^{i},\bm{\epsilon}^{i})),\bm{c})$, where $\bm{x}_{0}^{i}$ is sampled with SDE and $\text{ODE}^{k-1}(\text{SDE}(\bm{x}_{k}^{i},\bm{\epsilon}^{i}))$ is sampled with ODE-SDE-ODE. Trajectory branching allows for the precise attribution of the terminal reward to step $k$, effectively creating a temporally-aware reward signal.
Figure 6: (Left) Strong correlation between reward standard deviation and noise level across generative timesteps. (Right) Scale term analysis reveals a fundamental mismatch in standard GRPO: scale terms are inversely proportional to noise levels, causing low-noise refinement steps to dominate optimization despite minimal impact on image content.
@@ 4.1.2 Noise-Aware Policy Weighting
Reweighting by Noise Level. We wonder if directly leverage the level of the exploration space itself as a proxy for a reweighting factor. We visualized the noise level alongside the reward standard deviation (Figure 6, left) and observed a striking correspondence between the two. This strong correlation suggests that the noise level, serves as an excellent and intrinsic proxy for the exploration capacity and associated risk at each timestep.
@@ 4.2 Policy Gradient-Based Theoretical Justification
Consider a generative process parameterized by $\theta$, the policy gradient can be written as:
This yields:
This reveals that the natural gradient coefficient is proportional to $\sqrt{\frac{\Delta k(1-k)}{k}}$, which captures the intrinsic exploration potential at timestep $k$. After reweighting, we have the following derivation:
[…]
4.3 Group Strategy
Batch Group & Prompt Group. The original GRPO Shao et al. (2024) framework employs a group-level strategy, where trajectories are grouped based on a shared prompt. To mitigate issues of overfitting and instability, Reinforce++ Hu et al. (2025) later introduced a batch-level normalization.
Seed Group. In our work, trajectory branching requires $K$ distinct explorations at each timestep. Furthermore, we propose a novel seed-level grouping strategy. Under this approach, trajectories originating from the same prompt are further grouped if they share an identical initial noise seed. This methodology effectively controls for the influence of the initial noise, thereby ensuring that variations in reward can be attributed solely to the exploration conducted during the branching process. The experimental results, presented in Figure 7, validate our approach. As shown, TempFlow-GRPO consistently achieves superior performance compared to Flow-GRPO, regardless of the grouping strategy employed.
5 Experiment
Figure 7: Performance of TempFlow-GRPO and Flow-GRPO on different group strategy.
Following Flow-GRPO, we validate our approach on the Compositional Image Generation benchmark in Geneval Ghosh et al. (2023) and the Human Preference Alignment task in PickScore Kirstain et al. (2023). To ensure a fair comparison, our experimental setup is kept consistent with that of Flow-GRPO. For instance, we normalized the weights applied to the policy loss to have a mean of 1 at all timesteps. While our method introduces a trajectory branching strategy, we maintain an identical group configuration by setting 4 initial noise seeds and a branching factor ($K$) of 6. This results in a total group size of 24 (4 seeds $\times$ 6 branches) and 48 groups, perfectly matching the setup in Flow-GRPO. Furthermore, to benchmark our method against concurrent work, we conduct additional experiments on the FLUX.1-dev model. More detailed experimental settings (include comparsion with DanceGRPO) can be found in the Appendix A.2.
5.1 Main Results
Model Step Overall $\uparrow$ Single Obj. $\uparrow$ Two Obj. $\uparrow$ Counting $\uparrow$ Colors $\uparrow$ Position $\uparrow$ Attr. Binding $\uparrow$
Diffusion Models
LDM Rombach et al. (2022) - 0.37 0.92 0.29 0.23 0.70 0.02 0.05
SD1.5 Rombach et al. (2022) - 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD2.1 Rombach et al. (2022) - 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL Podell et al. (2023) - 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 Ramesh et al. (2022) - 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 Betker et al. (2023) - 0.67 0.96 0.87 0.47 0.83 0.43 0.45
Autoregressive Models
Show-o Xie et al. (2024b) - 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Emu3-Gen Wang et al. (2024) - 0.54 0.98 0.71 0.34 0.81 0.17 0.21
JanusFlow Ma et al. (2025a) - 0.63 0.97 0.59 0.45 0.83 0.53 0.42
Janus-Pro-7B Chen et al. (2025) - 0.80 0.99 0.89 0.59 0.90 0.79 0.66
GPT-4o Hurst et al. (2024) - 0.84 0.99 0.92 0.85 0.92 0.75 0.61
Flow Matching Models
FLUX.1 Dev Black et al. (2025) - 0.66 0.98 0.81 0.74 0.79 0.22 0.45
SD3.5-L Esser et al. (2024) - 0.71 0.98 0.89 0.73 0.83 0.34 0.47
SANA-1.5 4.8B Xie et al. (2025) - 0.81 0.99 0.93 0.86 0.84 0.59 0.65
SD3.5-M Esser et al. (2024) - 0.63 0.98 0.78 0.50 0.81 0.24 0.52
GRPO based Methods
SD3.5-M+Flow-GRPO Liu et al. (2025) 3800 0.88 0.99 0.96 0.90 0.87 0.83 0.78
SD3.5-M+TempFlow-GRPO 3800 0.97 1.00 1.00 0.96 0.95 0.99 0.91
Compositional Image Generation. We evaluate the compositional image generation capability of TempFlow-GRPO on the Geneval benchmark with its corresponding reward model. The experimental results are summarized in Table 1. As shown, our approach significantly improves the performance of the base model, increasing the overall score from 0.63 to 0.97. Furthermore, among GRPO-based methods, our method substantially outperforms Flow-GRPO: it achieves a performance of 0.97 within only 3,800 steps, whereas Flow-GRPO reaches only 0.88 under the same conditions. Additionally, as illustrated in Figure 4, our method requires only about 2,000 steps to achieve a score of 0.95, while Flow-GRPO needs approximately 5,600 steps to reach the same level. Overall, these results demonstrate that TempFlow-GRPO not only accelerates convergence but also achieves superior final performance compared to existing approaches.
Human Preference Alignment. To further validate the generalizability of our approach, we conducted experiments on the PickScore benchmark, using PickScore as the reward model. As shown in Figure 4 (left), TempFlow-GRPO, achieves the highest performance, surpassing the original Flow-GRPO by approximately 1.7% and outperforming the improved baseline Flow-GRPO (Prompt) by about 1.0%. Notably, our method requires only 100–200 training steps to match the performance of Flow-GRPO, and just 300–400 steps to reach the level of Flow-GRPO (Prompt). These results on PickScore further demonstrate the general applicability of our method as a unified flow-based RL algorithm across different reward models.
Figure 8: The performance and KL loss (smoothed) of HPSv3 on PickScore dataset (FLUX.1-dev).
HPSv3 on FLUX.1-dev. To further validate the performance of TempFlow-GRPO across different model and reward models, we select FLUX.1-dev as our base model with 1024 resolution and employ HPSv3 as the reward model. As shown in Figure 8, which presents the results on PickScore, our method requires only 80 steps to match the performance that Flow-GPO achieves in 300 steps. Moreover, our approach maintains a lower and more stable KL loss. Further experimental results and visualizations for the FLUX model can be found in the Appendix. A.3.
5.2 Analysis
Figure 9: Ablation studies on trajectory branch and noise-aware policy reweighting.
[…]
Ablation of TempFlow-GRPO. We conducted ablation studies to explore the effectiveness of our proposed components: the trajectory branch, noise-aware policy weighting and seed group. These ablations were performed on both the Geneval and PickScore benchmarks. As shown in the second and third subfigures of Figure 9, on the PickScore benchmark, introducing the trajectory branch further improves the performance of Flow-GRPO (Prompt), and applying noise-aware reweighting on top of this further boosted the performance. Finally, we achieved the highest performance by applying seed group. On the Geneval benchmark, the benefit of the noise-aware strategy is even more significant: compared to Flow-GRPO, noise-aware policy reweighting boosts performance from 0.82 to 0.92 in 1200 step, a 10% improvement while the trajectory branch also brings about a substantial gain of approximately 5% and seed group achieves 2% improvement. These ablation results clearly demonstrate the effectiveness of our proposed methods.
Qualitative Result. We also conducted qualitative analyses on the FLUX.1-dev, Flow-GRPO (Prompt), and TempFlow-GRPO. As shown in Figure10, compared to Flow-GRPO (Prompt), TempFlow-GRPO produces images with noticeably finer details and fewer visual artifacts or mistakes. In particular, our approach demonstrates superior capability in preserving complex structures and realistic textures. These qualitative improvements further highlighting the advantages of our method in generating high-quality, visually appealing images.
Figure 10: Qualitative comparison between FLUX.1-dev, Flow-GRPO (Prompt) and TempFlow-GRPO with HPSv3 as reward on PickScore prompts.
6 Conclusion
We presented TempFlow-GRPO, a temporally-aware reinforcement learning framework that addresses fundamental limitations in existing flow-based GRPO methods. Through trajectory branching, we enable precise credit assignment to intermediate actions without requiring specialized process reward models. Through noise-aware weighting, we ensure that optimization intensity matches each timestep’s exploration potential. We also introduce a principled seed group strategy that controls for initialization effects. Our extensive experiments demonstrate that TempFlow-GRPO achieves SOTA performance on human preference.
Limitations. Although our method achieves significant improvements in both performance and image quality, the current experiments focus primarily on algorithmic innovations rather than reward model enhancements. In future work, we plan to explore how to effectively incorporate multi-modal rewards from more powerful foundation models and develop comprehensive reward frameworks, aiming to enhance performance across multiple evaluation dimensions.
We utilize LLMs to assist with formula derivations and writing refinement on this paper.
Appendix A Appendix
A.1 Policy Gradient-Based Theoretical Framework